OpenStudy (anonymous):

The volume for a rectangular prism is given by the formula V = l · w · h, where l is the length of the prism, w is the width of the prism, and h is the height of the prism. The volume of a rectangular prism is 729 cubic inches. If the length and height are 9 inches each, what is the width of the rectangular prism? 81 inches 711 inches 72 inches 9 inches

OpenStudy (anonymous):

@nirmalnema @Nnesha @sammixboo

OpenStudy (unklerhaukus):

\[V = l \cdot w \cdot h\] rearrange the equation to solve for \(w\) the width of the prism

OpenStudy (anonymous):

81

sammixboo (sammixboo):

\(\rm V~=~l \times w \times h\) Since we know the Length and Height are 9, and the Volume is 729, we plug in the information \(\rm V~=~l \times w \times h\) \(\rm 729~=~9 \times w \times 9\) Now we multiply 9 and 9 So what is 9 x 9?

sammixboo (sammixboo):

The Answer to the problem isn't 81

sammixboo (sammixboo):

So what is 9 x 9?

OpenStudy (anonymous):

81

sammixboo (sammixboo):

Good now we have \(\rm 729~=~81\times w \) To solve this we must divide 81 on both sides of the equation \(\rm 729/81~=~81/81 \times w \) \(\rm 729/81~=~w \) So what is 729/81?

OpenStudy (nirmalnema):

w=729/81=9

sammixboo (sammixboo):

:( PLease let the asker solve this =) Thanks <3

OpenStudy (anonymous):

9?

sammixboo (sammixboo):

But it is 9, so that means 9 = w

sammixboo (sammixboo):

So 9 is our final answer

OpenStudy (anonymous):

thanks you both

sammixboo (sammixboo):

No problem =)